# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ 5.2 Main Results
Table 1: Comparison of performance under different dynamics shift scenarios and dataset types after 40K environment interactions. MR$=$Medium Replay, M $=$ Medium, ME $=$ Medium Expert. We mark a cell green if the method achieves the highest average return and improves over the second-best by at least 2%. Yellow highlights indicate practical ties (within 2% of the top method).
@@
(1) CompFlow consistently outperforms recent off-dynamics RL baselines across varying dataset qualities and types of dynamics shifts. Specifically, it achieves the highest return on 10 out of 15 tasks and ties on 4. On average, CompFlow attains a score of 2036, while the second-best baseline, H2O, achieves only 1710.
(2) CompFlow significantly outperforms the base algorithm BC-SAC in most scenarios. Notably, CompFlow achieves higher returns than BC-SAC on 12 out of 15 tasks under both morphology and kinematic shifts, and achieves ties on the remaining 3 tasks. On average, CompFlow yields an improvement of 20.1% over BC-SAC, clearly validating the effectiveness of our method.
(3) Additionally, we observe that recent baselines often perform similarly to BC-SAC across many tasks, consistent with the findings in [33], suggesting that they fail to effectively leverage the offline source domain data. This may stem from methods like H2O and BC-PAR relying on KL divergence or mutual information, which can be ill-defined under large dynamics gaps or disjoint supports. Meanwhile, BC-VGDF adopts a data filtering strategy based on estimated value functions, which is often more challenging than learning transition dynamics due to bootstrapping bias and target non-stationarity. In contrast, our method, CompFlow, leverages flow matching to model complex transition dynamics and exploits its theoretical connection to optimal transport to estimate the dynamics gap via Wasserstein distance—a more robust and principled metric.
5.3 Ablation and Hyperparameter Analysis
@@
Impact of data selection ratio $\xi\%$.
The data selection ratio $\xi$ decides how many source domain data in a sampled batch can be shared for policy training. A larger $\xi$ indicates that more source domain data will be admitted. To examine its influence, We sweep $\xi$ across {70, 50, 30, 20}. The results is shown in Figure 4. Although the optimal $\xi$ seems task dependent, moderate values of $\xi$ (e.g., 30 or 50) generally yield good performance across tasks, striking a balance between leveraging useful source data and avoiding high dynamics mismatch. When $\xi$ is too large (e.g., 70), performance often degrades—particularly in Walker Medium (Morph) and Walker Medium Replay (Morph)—likely due to the inclusion of low-quality transitions with large dynamics gap. HalfCheetah consistently achieves higher returns compared to other domains, suggesting that knowledge transfer and policy learning in this environment are easier. Consequently, overly conservative filtering (e.g., $\xi=20$) may exclude valuable source data, leading to slower learning. In this case, allowing more source data (e.g., $\xi=70$) appears beneficial.
@@ 6 Conclusion and Limitations
In this paper, we proposed CompFlow, a new method for estimating the dynamics gap in reinforcement learning with shifted-dynamics data. CompFlow leverages the theoretical connection between flow matching and optimal transport. To address data scarcity in the target domain, we adopt a composite flow structure that builds the target flow model on top of the output distribution from the source domain flow. This composite formulation improves generalization and enables the use of Wasserstein distance between source and target transitions as a robust measure of the dynamics gap. Using this pricinpled estimation, we further encourage the policy to explore regions with high dynamics gap and provide a theoretical analysis of the benefits. Empirically, we demonstrate that CompFlow consistently outperforms or matches state-of-the-art baselines across 15 diverse RL tasks with varying types of dynamics shift.
Limitations. (1) CompFlow is currently limited to settings where the source and target domains share the same state and action spaces. (2) Our evaluation is confined to simulated environments; evaluating CompFlow in real-world scenarios remains an important direction for future work. (3) Extending CompFlow to multi-domain transfer settings is also a promising avenue for exploration.
@@
[…]
Appendix A Broader Impacts
Our work aims to make reinforcement learning more practical in real-world domains—such as healthcare, robotics, and conservation—where online interaction is costly or unsafe. By addressing dynamics shifts between offline and online environments, our method enables more reliable and sample-efficient policy learning. This can support safer deployment in high-stakes settings, such as clinical decision support or adaptive anti-poaching strategies. However, we emphasize that policies trained on historical data should be used with caution, as misaligned dynamics or biased datasets may lead to unintended outcomes if not properly validated.
Appendix B Algorithms of Training Optimal Transport Flow Matching
@@
1:Dataset $D$, batch size $k$, cost function $c(\cdot,\cdot)$, learning rate $\mathsf{lr}$, flow model $v_{\theta}$
@@
4:     Sample data batch $\{x_{1}^{(j)}\}_{j=1}^{k}\sim D$
@@
11:         Compute target velocity:
$v^{*(\ell)}\doteq x_{1}^{(\ell)}-x_{0}^{(\ell)}$
12:     end for
13:     Compute loss:
$\mathcal{L}\doteq\frac{1}{k}\sum_{\ell=1}^{k}\left\|v_{\theta}(x_{t}^{(\ell)},t^{(\ell)})-v^{*(\ell)}\right\|_{2}^{2}$
14:     Update parameters:
@@
15:end while
16:return $\theta$
Appendix C Algorithms of Training Source and Target Flows
The full training algorithms of the source flow and target flow are given in Algorithm 2 and Algorithm 3 respectively.
Algorithm 2 Training of Source Domain Flow Matching
1:Source domain dataset $D_{\text{src}}$, batch size $k$, flow model $v_{\theta}$, learning rate $\mathsf{lr}$
@@ Appendix C Algorithms of Training Source and Target Flows
3:     Sample a batch of transitions $\{(s^{(i)},a^{(i)},s^{\prime(i)})\}_{i=1}^{k}\sim D_{\text{src}}$
4:     For each $i$, define target sample $x_{1}^{(i)}\doteq(s^{(i)},a^{(i)},s^{\prime(i)})$
@@
9:         Compute target velocity: $v^{\ast}_{i}=x_{1}^{(i)}-x_{0}^{(i)}$
10:     end for
11:     Compute training loss:
$\mathcal{L}=\frac{1}{k}\sum_{i=1}^{k}\left\|v_{\theta}(x_{t}^{(i)},t^{(i)},s^{(i)},a^{(i)})-v^{\ast}_{i}\right\|_{2}^{2}$
12:     Update $\theta\leftarrow\theta-\mathsf{lr}\nabla_{\theta}\mathcal{L}$
13:end while
14:return Trained parameters $\theta$
Algorithm 3 Training of Target Domain Flow via Optimal Transport
1:Pre-trained source flow model $\psi_{\theta}(x,t\mid s,a)$, source dataset $D_{\text{src}}$, target dataset $D_{\text{tar}}$, batch size $k$, cost function $c(\cdot)$, regularization weight $\eta$, learning rate $\mathsf{lr}$, target flow model $v_{\phi}$
@@
4:     Sample source batch $\{s_{\rm src}^{(i)},a_{\rm src}^{(i)}\}_{i=1}^{k}\sim D_{\rm src}$
@@
6:         Predict source next state:
$s^{\prime(i)}_{\rm src}\doteq x_{1}^{(i)}\doteq\psi_{\theta}(x_{0}^{(i)},1\mid s_{\rm src}^{(i)},a_{\rm src}^{(i)})$
@@
8:     Sample target batch $\{s_{\rm tar}^{(j)},a_{\rm tar}^{(j)},s^{\prime(j)}_{\rm tar}\}_{j=1}^{k}\sim D_{\rm tar}$
9:     Define $x_{2}^{(j)}\doteq s^{\prime(j)}_{\rm tar}$
@@
$c(i,j)\doteq\|x_{1}^{(i)}-x_{2}^{(j)}\|_{2}^{2}+\eta\left(\|s_{\rm src}^{(i)}-s_{\rm tar}^{(j)}\|^{2}_{2}+\|a_{\rm src}^{(i)}-a_{\rm tar}^{(j)}\|^{2}_{2}\right)$
@@
13:     Sample interpolation times $\{t^{(\ell)}\}_{\ell=1}^{k}\sim\mathcal{U}[1,2]$triangleright Time $t$ for the target flow is from 1 to 2
@@
$\mathcal{L}\doteq\frac{1}{k}\sum_{\ell=1}^{k}\left\|v_{\phi}(x_{t}^{(\ell)},t^{(\ell)},s_{t}^{(\ell)},a_{t}^{(\ell)})-v^{*(\ell)}\right\|_{2}^{2}$
@@
Appendix D Algorithm of CompFlow built on Soft-Actor-Critic
@@
1:Input: Source dataset $D_{\rm src}$, target environment $\mathcal{M}_{\rm tar}$, max interaction steps $T_{\rm max}$, target model training frequency $\mathsf{train\_freq}$, source data selection ratio $\xi$, gap reward scale $\beta$, batch size $B$, behavior cloning weight $\omega$, warmup steps $\mathsf{warmup\_steps}$, learning rate $\mathsf{lr}$, target update rate $\varpi$
2:Initialization: Policy $\pi_{\varphi}$, Q-functions $\{Q_{\varsigma_{i}}\}_{i=1,2}$, target Q-functions $\{Q_{\varsigma_{i}}^{\rm tgt}\}_{i=1,2}\leftarrow\{Q_{\varsigma_{i}}\}_{i=1,2}$, target replay buffer $D_{\rm tar}$
3:Pretrain source flow model $\psi_{\theta}(x,t\mid s,a)$ on $D_{\rm src}$ via Algorithm 2
@@
5:     Sample transition $(s_{\rm tar},a_{\rm tar},r_{\rm tar},s^{\prime}_{\rm tar})\sim\mathcal{M}_{\rm tar}$ using policy $\pi_{\varphi}$
6:     Update replay buffer: $D_{\rm tar}\leftarrow D_{\rm tar}\cup\{(s_{\rm tar},a_{\rm tar},r_{\rm tar},s^{\prime}_{\rm tar})\}$
@@
8:         Train target flow $\psi_{\phi}(x,t\mid s,a)$ on $D_{\rm tar}$ via Algorithm 3
@@
11:         Sample $b_{\rm src}=\{(s_{\rm src}^{(i)},a_{\rm src}^{(i)},r_{\rm src}^{(i)},s_{\rm src}^{\prime(i)})\}_{i=1}^{B}\sim D_{\rm src}$
12:         Sample $b_{\rm tar}=\{(s_{\rm tar}^{(i)},a_{\rm tar}^{(i)},r_{\rm tar}^{(i)},s_{\rm tar}^{\prime(i)})\}_{i=1}^{B}\sim D_{\rm tar}$
13:         Estimate $\widehat{\Delta}(s_{\rm src}^{(i)},a_{\rm src}^{(i)})$ for each $(s_{\rm src}^{(i)},a_{\rm src}^{(i)})\in b_{\rm src}$ via Eq. 2
14:         Select top $\xi\%$ of $b_{\rm src}$ with lowest gap: $\tilde{b}_{\rm src}$
15:         For each $(s,a,r)\in\tilde{b}_{\rm src}$, update reward: $r\leftarrow r+\beta\widehat{\Delta}(s,a)$
@@
$\nabla_{\varsigma_{i}}\mathcal{L}_{Q}\leftarrow$ $\frac{1}{B}\sum_{(s^{\rm tar},a^{\rm tar},s^{\prime\rm tar},r^{\rm tar})\in b_{\rm tar}}\nabla_{\varsigma_{i}}\left(Q_{\varsigma_{i}}(s^{\rm tar},a^{\rm tar})-y\right)^{2}$
$+\frac{1}{|\tilde{b}_{\rm src}|}\sum_{(s^{\rm src},a^{\rm src},s^{\prime\rm src},r^{\rm src})\in\tilde{b}_{\rm src}}\nabla_{\varsigma_{i}}\left(Q_{\varsigma_{i}}(s^{\rm src},a^{\rm src})-y\right)^{2}$
@@
$\lambda=\omega\left/\left\{\frac{1}{2B}\sum_{\tilde{s}\in\tilde{b}_{\rm src}\cup b_{\rm tar}}\left|\min\left\{Q_{\varsigma_{1}}(\tilde{s},\tilde{a}),Q_{\varsigma_{2}}(\tilde{s},\tilde{a})\right\}\right|\right\}\right.$
@@
$\nabla_{\varphi}\mathcal{L}_{\pi}\leftarrow$ $-\frac{\lambda}{|\tilde{b}_{\rm src}\cup b_{\rm tar}|}\sum_{\tilde{s}\in\tilde{b}_{\rm src}\cup b_{\rm tar}}\nabla_{\varphi}\left[\min_{i=1,2}Q_{\varsigma_{i}}(\tilde{s},\tilde{a})+\alpha\mathcal{H}[\pi_{\varphi}(\cdot|\tilde{s})]\right]$
$+\frac{1}{|b_{\rm src}|}\sum_{(s,a)\in b_{\rm src}}\nabla_{\varphi}\left\|a-\bar{a}\right\|^{2}$
@@
Appendix E Proof of Lemma 2.2
@@
Proof of Lemma  2.2 is already provided as an intermediate step for Theorem A.4 in [32], the offline performance bound case. We include their proof here with slight modifications for completeness.
We first cite the following two necessary lemmas also used in [32]. ∎
Lemma E.1 (Extended telescoping lemma).
@@
Proof. This is Lemma C.2 in [59]. $\square$
Lemma E.2.
@@
Proof. This is Lemma B.3 in [32]. $\square$

### v3

@@ 5.1.2 Main Results
Table 1: Comparison of return under different dynamics shift scenarios and dataset types after 40K environment interactions. MR = Medium Replay, M = Medium, ME = Medium Expert. A cell is green if the method has the highest mean and improves over the second best by at least 2%. Cells within 2% of the top mean are marked in yellow.
@@
(1) CompFlow consistently outperforms recent off-dynamics reinforcement learning baselines across a wide range of dataset qualities and types of dynamics shifts. Specifically, it achieves the highest return on 19 out of 27 tasks and matches the top performance on an additional 5 tasks. On average, CompFlow attains a score of 2193, compared to 1920 achieved by the second-best baseline, BC-SAC, representing a relative improvement of 14.2%. To further validate the robustness of our method, we present additional experimental results under even more pronounced dynamics shifts in Appendix J, which demonstrates consistent findings.
(2) Incorporating offline data can significantly enhance learning performance. On average, CompFlow achieves a 149.8% improvement over SAC. In addition, CompFlow outperforms the base algorithm BC-SAC, which directly incorporates all offline data, on 23 out of 27 tasks and matches its performance on the remaining 3, providing strong empirical support for the effectiveness of our method.
(3) Additionally, we observe that recent baselines often perform similarly to BC-SAC across many tasks, consistent with the findings in [lyu2024odrl], suggesting that they fail to effectively leverage the offline source domain data. This may stem from methods like H2O and BC-PAR relying on KL divergence or mutual information, which can be ill-defined under large dynamics gaps or different supports. Meanwhile, BC-VGDF adopts a data filtering strategy based on estimated value functions, which is often more challenging than learning transition dynamics due to bootstrapping bias and target non-stationarity. In contrast, our method, CompFlow, leverages flow matching to model complex transition dynamics and exploits its theoretical connection to optimal transport to estimate the dynamics gap via Wasserstein distance—a more robust and principled metric.
5.1.3 Ablation and Hyperparameter Analysis
@@
Impact of data selection ratio $\xi\%$. The data selection ratio $\xi$ decides how many source domain data in a sampled batch can be shared for policy training. A larger $\xi$ indicates that more source domain data will be admitted. To examine its influence, We sweep $\xi$ across {70, 50, 30, 20}. The results is shown in Figure 4. Although the optimal $\xi$ seems task dependent, moderate values of $\xi$ (e.g., 30 or 50) generally yield good performance across tasks, striking a balance between leveraging useful source data and avoiding high dynamics mismatch. When $\xi$ is too large (e.g., 70), performance often degrades, particularly in Walker Medium (Morph) and Walker Medium Replay (Morph), likely due to the inclusion of low-quality transitions with large dynamics gap. HalfCheetah consistently achieves higher returns compared to other domains, suggesting that knowledge transfer and policy learning in this environment are easier. Consequently, overly conservative filtering (e.g., $\xi=20$) may exclude valuable source data, leading to slower learning. In this case, allowing more source data (e.g., $\xi=70$) appears beneficial.
@@
5.2 Patrol Policy Learning for Wildlife Conservation
We evaluate CompFlow in a wildlife conservation setting, where the objective is to learn a patrol policy that allocates limited ranger effort to reduce poaching and protect wildlife. We adopt the simulation environment introduced by xu2021robust. The environment is represented as a spatial grid, partitioned into $1\times 1$ km cells. At each time step, the agent selects a distribution of patrol effort across the cells, subject to a fixed budget. The state, which includes wildlife population density and poaching risk for each cell, evolves over time according to dynamics that depend on both poacher behavior and the deployed patrols. The per-step reward reflects the expected number of animals preserved through deterrence and spatial coverage achieved by the patrol strategy.
Figure 6: Reward on the wildlife conservation task.
We assume access to an offline dataset collected under a previous policy in Murchison Falls National Park. The objective is to use this data to learn an improved patrol policy for Queen Elizabeth National Park, with only limited interaction with the new environment. Due to differences in ecological conditions and poacher behavior, the transition dynamics in the two parks are not the same. Figure 6 presents the results. CompFlow achieves the highest reward, exceeding the best baseline, BC-PAR, by 8.8%. Compared to training a policy from scratch using SAC, CompFlow improves the average reward by 20.8% under the same interaction budget. This improvement is especially important in conservation settings where ranger capacity is limited and the protected area is large. For instance, Murchison Falls National Park spans approximately 3,900 square kilometers, while Queen Elizabeth National Park covers about 1,980 square kilometers. The large size of these areas makes full coverage difficult, and learning effective patrol strategies with minimal exploration is crucial for protecting wildlife.
@@ 6 Conclusion and Limitations
In this paper, we proposed CompFlow, a new method for estimating the dynamics gap in reinforcement learning with shifted-dynamics data. CompFlow leverages the theoretical connection between flow matching and optimal transport. To address data scarcity in the online environment, we adopt a composite flow structure that builds the online flow model on top of the output distribution from the offline flow. This composite formulation improves generalization and enables the use of Wasserstein distance between offline and online transitions as a robust measure of the dynamics gap. Using this pricinpled estimation, we further encourage the policy to explore regions with high dynamics gap and provide a theoretical analysis of the benefits. Empirically, we demonstrate that CompFlow consistently outperforms or matches state-of-the-art baselines across diverse RL tasks with varying types of dynamics shift.
Due to the space limit, we discuss the limitations of CompFlow in Appendix A.
Acknowledgement
We are thankful to the Uganda Wildlife Authority for granting us access to incident data from Murchison Falls and Queen Elizabeth National Park. We also thank the anonymous reviewers for their valuable feedback. This work was supported by ONR MURI N00014-24-1-2742.
@@ Acknowledgement
Appendix A Limitations
(1) CompFlow is currently limited to settings where the offline data and the online environment share the same state and action spaces. Future work could explore estimating the dynamics gap in a shared latent embedding space to relax this assumption [zhuang2023dygen]. (2) Our evaluation is conducted entirely in simulated environments; applying CompFlow to real-world scenarios remains an important direction for future work. (3) We have not yet incorporated the uncertainty in estimating the dynamics gap. Future work may explore how to quantify this uncertainty and use it to improve both data filtering and exploration strategies [kong2023uncertainty, li2024muben].
Appendix B Broader Impacts
Our work aims to make reinforcement learning more practical in real-world domains such as healthcare, robotics, and conservation, where online interaction is often costly, limited, or unsafe. By addressing the dynamics shift between offline data and the online environment, our method enables more reliable and sample-efficient policy learning. This capability supports safer deployment in high-stakes applications, including clinical decision support and adaptive anti-poaching strategies. Nonetheless, we emphasize that policies trained on historical data should be applied with caution, as misaligned dynamics or biased datasets may lead to unintended consequences if not carefully validated.
Appendix C Algorithms of Training Optimal Transport Flow Matching
@@
1:Dataset $\mathcal{D}$, batch size $k$, cost function $c(\cdot,\cdot)$, learning rate $\mathsf{lr}$, flow model $v_{\theta}$
@@
4:  Sample data batch $\{x_{1}^{(j)}\}_{j=1}^{k}\sim\mathcal{D}$
@@
11:  end for
12:  Compute loss:
$\mathcal{L}\doteq\frac{1}{k}\sum_{\ell=1}^{k}\left\|v_{\theta}(x_{t}^{(\ell)},t^{(\ell)})-(x_{1}^{(\ell)}-x_{0}^{(\ell)})\right\|_{2}^{2}$
13:  Update parameters:
@@
14:end while
15:return $\theta$
Appendix D Algorithms of Training Offline and Online Flows
The full training algorithms of the offline flow and online flow are given in Algorithm 2 and Algorithm 3 respectively.
Algorithm 2 Training of Offline Flow Matching
1:Offline dataset $\mathcal{D}_{\text{off}}$, batch size $k$, flow model $v_{\theta}$, learning rate $\mathsf{lr}$
@@ Appendix D Algorithms of Training Offline and Online Flows
3:  Sample a batch of transitions $\{(s^{(i)},a^{(i)},s^{\prime(i)})\}_{i=1}^{k}\sim\mathcal{D}_{\text{off}}$
4:  For each $i$, define target sample $x_{1}^{(i)}\doteq s^{\prime(i)}$
@@
9:  end for
10:  Compute training loss:
$\mathcal{L}=\frac{1}{k}\sum_{i=1}^{k}\left\|v_{\theta}(x_{t}^{(i)},t^{(i)},s^{(i)},a^{(i)})-(x_{1}^{(i)}-x_{0}^{(i)})\right\|_{2}^{2}$
11:  Update $\theta\leftarrow\theta-\mathsf{lr}\nabla_{\theta}\mathcal{L}$
12:end while
13:return Trained parameters $\theta$
Algorithm 3 Training of Online Flow via Optimal Transport
1:Pre-trained offline flow model $\psi_{\theta}(x,t\mid s,a)$, offline dataset $\mathcal{D}_{\text{off}}$, target dataset $\mathcal{D}_{\text{on}}$, batch size $k$, cost function $c(\cdot)$, regularization weight $\eta$, learning rate $\mathsf{lr}$, target flow model $v_{\phi}$
@@
4:  Sample offline batch $\{s_{\rm off}^{(i)},a_{\rm off}^{(i)}\}_{i=1}^{k}\sim\mathcal{D}_{\rm off}$
@@
6:   Predict next state for the offline environment:
$s^{\prime(i)}_{\rm off}\doteq x_{1}^{(i)}\doteq\psi_{\theta}(x_{0}^{(i)},1\mid s_{\rm off}^{(i)},a_{\rm off}^{(i)})$
@@
8:  Sample online batch $\{s_{\rm on}^{(j)},a_{\rm on}^{(j)},s^{\prime(j)}_{\rm on}\}_{j=1}^{k}\sim\mathcal{D}_{\rm on}$
9:  Define $x_{2}^{(j)}\doteq s^{\prime(j)}_{\rm on}$
@@
$c(i,j)\doteq\|x_{1}^{(i)}-x_{2}^{(j)}\|_{2}^{2}+\eta\left(\|s_{\rm off}^{(i)}-s_{\rm on}^{(j)}\|^{2}_{2}+\|a_{\rm off}^{(i)}-a_{\rm on}^{(j)}\|^{2}_{2}\right)$
@@
13:  Sample interpolation times $\{t^{(\ell)}\}_{\ell=1}^{k}\sim\mathcal{U}[1,2]$ $\triangleright$ Time $t$ for the online flow is from 1 to 2
@@
$\mathcal{L}\doteq\frac{1}{k}\sum_{\ell=1}^{k}\left\|v_{\phi}(x_{t}^{(\ell)},t^{(\ell)},s_{\rm on}^{(\ell)},a_{\rm on}^{(\ell)})-(x_{2}^{(j_{\ell})}-x_{1}^{(i_{\ell})})\right\|_{2}^{2}$
@@
Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
@@
1:Input: Offline dataset $\mathcal{D}_{\rm off}$, Online environment $\mathcal{M}_{\rm on}$, max interaction steps $T_{\rm max}$, target model training frequency $\mathsf{train\_freq}$, source data selection ratio $\xi$, gap reward scale $\beta$, batch size $B$, behavior cloning weight $\omega$, warmup steps $\mathsf{warmup\_steps}$, learning rate $\mathsf{lr}$, target update rate $\varpi$
2:Initialization: Policy $\pi_{\varphi}$, Q-functions $\{Q_{\varsigma_{i}}\}_{i=1,2}$, target Q-functions $\{Q_{\varsigma_{i}}^{\rm tgt}\}_{i=1,2}\leftarrow\{Q_{\varsigma_{i}}\}_{i=1,2}$, target replay buffer $\mathcal{D}_{\rm on}$
3:Pretrain offline flow model $\psi_{\theta}(x,t\mid s,a)$ on $\mathcal{D}_{\rm off}$ via Algorithm 2
@@
5:  Sample transition $(s_{\rm on},a_{\rm on},r_{\rm on},s^{\prime}_{\rm on})\sim\mathcal{M}_{\rm on}$ using policy $\pi_{\varphi}$
6:  Update replay buffer: $\mathcal{D}_{\rm on}\leftarrow\mathcal{D}_{\rm on}\cup\{(s_{\rm on},a_{\rm on},r_{\rm on},s^{\prime}_{\rm on})\}$
@@
8:   Train target flow $\psi_{\phi}(x,t\mid s,a)$ on $\mathcal{D}_{\rm on}$ via Algorithm 3
@@
11:   Sample $b_{\rm off}=\{(s_{\rm off}^{(i)},a_{\rm off}^{(i)},r_{\rm off}^{(i)},s_{\rm off}^{\prime(i)})\}_{i=1}^{B}\sim D_{\rm off}$
12:   Sample $b_{\rm on}=\{(s_{\rm on}^{(i)},a_{\rm on}^{(i)},r_{\rm on}^{(i)},s_{\rm on}^{\prime(i)})\}_{i=1}^{B}\sim\mathcal{D}_{\rm on}$
13:   Estimate $\widehat{\Delta}(s_{\rm off}^{(i)},a_{\rm off}^{(i)})$ for each $(s_{\rm off}^{(i)},a_{\rm off}^{(i)})\in b_{\rm off}$ via Eq. 3
14:   Select top $\xi\%$ of $b_{\rm off}$ with lowest gap: $\tilde{b}_{\rm off}$
15:   For each $(s,a,s^{\prime},r)\in\tilde{b}_{\rm off}$, update reward: $r\leftarrow r+\beta\widehat{\Delta}(s,a)$
@@
$\nabla_{\varsigma_{i}}\mathcal{L}_{Q}\leftarrow$ $\frac{1}{B}\sum_{(s^{\rm on},a_{\rm on},s^{\prime}_{\rm on},r_{\rm on})\in b_{\rm on}}\nabla_{\varsigma_{i}}\left(Q_{\varsigma_{i}}(s_{\rm on},a_{\rm on})-y\right)^{2}$
$+\frac{1}{|\tilde{b}_{\rm off}|}\sum_{(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off},r_{\rm off})\in\tilde{b}_{\rm off}}\nabla_{\varsigma_{i}}\left(Q_{\varsigma_{i}}(s_{\rm off},a_{\rm off})-y\right)^{2}$
@@
[…]
@@
[…]
$+\frac{1}{|b_{\rm off}|}\sum_{(s,a)\in b_{\rm off}}\nabla_{\varphi}\left\|a-\bar{a}\right\|^{2}$
@@
Appendix F Proof of Lemma 2.2
@@
Proof of Lemma  2.2 is already provided as an intermediate step for Theorem A.4 in [lyu2024cross], the offline performance bound case. We include their proof here with slight modifications for completeness.
We first cite the following two necessary lemmas also used in [lyu2024cross]. ∎
Lemma F.1 (Extended telescoping lemma).
@@
Proof. This is Lemma C.2 in [xu2023cross]. $\square$
Lemma F.2.
@@
Proof. This is Lemma B.3 in [lyu2024cross]. $\square$